OpenStudy (anonymous):

A strong rope needs to be fixed with one end atached to the top of a 5m pole and the other end pegged at an angle of 60 degrees with the level ground. Find the required length of the rope, correct to 2 decimal places.

OpenStudy (anonymous):

the answer is 5.77m i just dont know how to get it

OpenStudy (harsimran_hs4):



OpenStudy (harsimran_hs4):

now suppose rope has length x mts so, sin(60) = 5/x x = 5/sin(60) = 5/((sqrt 3)/2)